OpenStudy (anonymous):

1.Zach borrowed four thousand dollars from his grandfather to buy a used car so he could travel to and from work. After working for 2 years, he has saved five thousand two hundred dollars to pay back to his grandfather. At what interest rate did Zach borrow the money? Round to the nearest whole percent. twelve percent fifteen percent eighteen percent thirty-eight percent

OpenStudy (anonymous):

Do you use this formula I=prt

OpenStudy (johnweldon1993):

The simple interest formula yes...rearrange and solve for 'r'

OpenStudy (anonymous):

How?

OpenStudy (johnweldon1993):

Well...we know that the original simple interest formula is I = prt where I = interest earned P = principal (original amount) R = rate of interest and T = time (in years) Since we already have I, P, and T...we want R so instead of writing this as 1200 = 4000*R*2 we could rearrange the original equation...by dividing both sides by PT I = PRT ------- PT PT this will leave us with \[R = \frac{ I }{ PT }\] So fill in your values and what do you get...?

OpenStudy (johnweldon1993):

And just to make sure everything is clear I (interest earned) is what you get when you do (current balance - original balance)

OpenStudy (anonymous):

so wait you divide two times?

OpenStudy (johnweldon1993):

Not quite....you know what...forget what I said about isolating the variable for the moment...lets do it the other way You have the equation I = PRT right? and you have what I, P, and T are right? so just put in those values 1200 = 4000*R*2 what does R equal...?

OpenStudy (anonymous):

how did you get 1200

OpenStudy (johnweldon1993):

Remember when I said that I (which is interest earned) is found by taking the current balance (5200) and subtracting the original balance (4000)

OpenStudy (anonymous):

ok

OpenStudy (johnweldon1993):

So now that you have the I, P, and T....you can solve for R (rate)

OpenStudy (anonymous):

B

OpenStudy (anonymous):

right?

OpenStudy (johnweldon1993):

B) Fifteen Percent would be correct! Good job!

OpenStudy (anonymous):

Thank you

OpenStudy (johnweldon1993):

No problem!